OpenStudy (anonymous):

what is a particular solution for ODE and what is general solution for ODE? And for ODE, differential operator, if I got : [(D-6)^2+32]y=0, then how can I get y?

OpenStudy (anonymous):

i know (D-6)^2=e^6x D^2 e^-6x, but i dont know how to solve it when it plus a constant 32

OpenStudy (anonymous):

@TuringTest Hi, could you help we solve this?

OpenStudy (anonymous):

D=d/dx

OpenStudy (anonymous):

?

OpenStudy (anonymous):

yes